OpenStudy (anonymous):

For my intermediate algebra class, I need to find the distance between the given points: (√3, 5√2) and (√3, √2) The formula we have been given was: d=√(x2-x1)^2 + (y2-y1)^2 I have an idea of what to do but I'm not sure if it's right.

OpenStudy (anonymous):

here x1=root3 y1=5.root3 x2=root3 y2=root2 apply these values in your formula you will get the answer

OpenStudy (anonymous):

I know that much, but I get stuck after I plug in the numbers. Are the square roots inside the square root gone?

OpenStudy (anonymous):

\[(\sqrt{3}-\sqrt{3})^2+(\sqrt{2}-5\sqrt{2})^2\] \[(4\sqrt{2})^2\] \[4\times4\times \sqrt{2}\times \sqrt{2}\] \[16\times2=32\]

OpenStudy (anonymous):

Is it okay to add square roots? Or are they their own separate entities?

OpenStudy (anonymous):

yes see roots as the variables x=root2 then x-5x=?

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (anonymous):

please identify the best response to help others